OpenStudy (anonymous):

|sin|x| | = | |X| - 1 | how many solutions of this equation is possible ?

OpenStudy (anonymous):

@genius12 nd @satellite73 .. please help. will give medal . :P

OpenStudy (primeralph):

is x = X?

OpenStudy (anonymous):

ya ya .. ! i mistakenly wrote dat .

OpenStudy (anonymous):

oh god .. !! my head is steaming.. ! please help.!

OpenStudy (anonymous):

Note that the left-hand side is the same as saying \(\bf |sin(x)|\) It should also be obvious that there will be 4 solutions. But we only need to find 2 solutions since the other two are just the "negative versions of the positive solutions. Two find these 2 positive solutions, we can think of finding the solution to the following:\[\bf \sin(x)=-x+1\]and\[\bf \sin(x)=x-1\]Both cases yield 1 solution each for a total of 2 solutions.

OpenStudy (anonymous):

@tanjeetsarkar96

OpenStudy (anonymous):

remove mode from x (both x) then

OpenStudy (anonymous):

|sinx| = |x-1|

OpenStudy (anonymous):

how is it possible ??? @McLove

OpenStudy (anonymous):

@McLove, you can't do that.\[\bf |x-1| \ne ||x|-1|\]Like I said, you can only do it for the left-side of the equation.

OpenStudy (anonymous):

oh sorry i didn't saw negative sign lol

OpenStudy (anonymous):

@tanjeetsarkar96 I have simplified the steps for you but ideally, for equations like these, you are bound to use graphing calculators.

OpenStudy (anonymous):

Because finding the intersection between a line and a sine function cannot be done normally.

OpenStudy (anonymous):

the ques is sin |x| not sinx ..!

OpenStudy (anonymous):

I know lol. But like I said:\[\bf |\sin|x||=|\sin(x)|\] @tanjeetsarkar96

OpenStudy (anonymous):

So you can remove the absolute value around x

OpenStudy (anonymous):

Wait. You don't actually need the solutions just how many there are? I didn't realise that lol. Like I stated in my first post, there will be 4 solutions. @tanjeetsarkar96

OpenStudy (anonymous):

i still dont get u @genius12

OpenStudy (anonymous):

If you graph the left handside of the equation and the right hand side, you get this:4 intersection points. You don't need a graphing calculator to graph these. It should be straight-forward. @tanjeetsarkar96